OpenStudy (anonymous):

The Congress of the United States signed the Declaration of Independence in 1776. Find the difference in years between the year that the Congress signed this document and the year that you were born.

OpenStudy (anonymous):

Take whatever year you were born in and subtract 1776.

OpenStudy (anonymous):

Ok what year were you born in?

OpenStudy (anonymous):

it need to be in an equation form

OpenStudy (anonymous):

let x = year you were born let y = year congress singed the declaration

OpenStudy (anonymous):

y=x-1776. Where y is the difference and x is the year you were born.

OpenStudy (anonymous):

would this be right x=1776+ 2001

OpenStudy (anonymous):

or halo's way is actually better

OpenStudy (anonymous):

No, you'd be adding 1776 to 2001. That is definitely not the difference.

OpenStudy (anonymous):

thats why i'm confused because then i have to solve it by doing the same on both sides.

OpenStudy (anonymous):

I was born in 1997. 1997-1776=221 So there is 221 years between the year the constitution was signed and the year I was born. Does that make sense?

OpenStudy (anonymous):

yea it does but not when i have to write out the equation and then solve it

OpenStudy (anonymous):

if i wrote y= 2001 -1776 for the equation then i write it out then under it i have to do the oppiste it adding it doesnt work

OpenStudy (anonymous):

right if its subtracation i need to do the oppiste addition to balance out the sides. Its asking for the difference which is to subtract.